During a semester, a class of n=50 students took two exams: a midterm and final. The mean source for the midterm exam is 80 points with the standard deviation of 6 points. The mean score for the final exam is 85 points with the standard deviation of 10 points. The calculated coefficient of correlation (r) for the two scores is 0.75.

Q. Calculate the value of SSE (sum of square of error) using the above information.
Answer #1

here SSE =(1-r)2*SST

from above we get SST =(n-1)sy2 =(50-1)*102 =4900

therefore SSE =(1-0.752)*4900=2143.75